OpenStudy (anonymous):

A ball is thrown into the air with an upward velocity of 36 ft/s. Its height h in feet after t seconds is given by the function . In how many seconds does the ball reach its maximum height? Round to the nearest hundredth if necessary. What is the ball’s maximum height? 2.25 s; 9 ft 1.13 s; 31.5 ft 1.13 s; 69.75 ft 1.13 s; 29.25 ft

OpenStudy (ivettef365):

what is the function?

OpenStudy (anonymous):

h=-16^2+36+9

OpenStudy (ivettef365):

first use formula -b / 2a so you have -36/2(-16) = 36/32 = 9/8 = 1.13

OpenStudy (ivettef365):

then do f(9/8) =16t^2 + 36t + 9

OpenStudy (ivettef365):

f(9/8) = 69.75 confirm if that is what you got.

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

ok :)